OpenStudy (anonymous):

what is cot x when sin x is 1/2

OpenStudy (anonymous):

you have no way of knowing unless you know what quadrant you are in

OpenStudy (anonymous):

tan is square root of 3 over 3 cos is - square root of 3over 2

OpenStudy (anonymous):

\[\sin(x)=\frac{1}{2}\] \[\cos(x)=\pm\frac{\sqrt{3}}{2}\]

OpenStudy (lgbasallote):

sin x = O/H A = sqrt H^2 - O^2 A = sqrt 3 cot = A/O so sqrt 3/1 which is simply sqrt 3

OpenStudy (anonymous):

well it might be \[\sqrt{3}\] then again it might be \[-\sqrt{3}\]

OpenStudy (lgbasallote):

yeah... plus or minus